Peter is twice as old as Sam. If 8 is subtracted from Sam’s age and 4 is added to Peter’s age,
Peter will then be four times as old as Sam. Find their ages.

Answer by ikleyn(52803)

.
Peter is twice as old as Sam. If 8 is subtracted from Sam’s age and 4 is added to Peter’s age,
Peter will then be four times as old as Sam. Find their ages.
~~~~~~~~~~~~~~~~~

Sam   =  x years old.
Peter = 2x years old.


Equation

    8 is subtracted from Sam’s age        4 is added to Peter’s age
          4 * (x-8)                  =    (2x+4)


Simplify and find x

    4x - 32 = 2x + 4

    4x - 2x = 32 + 4

       2x   =   36

        x   =   36/2 = 18


ANSWER.  Sam is 18 years old.  Peter is 2*18 = 36 years old.

Solved.